OpenStudy (mathhelp346):

The net price of a refrigerator is $455 when a 30% discount is allowed. What is its list price? I got $318.50. Is this correct?

OpenStudy (gw2011):

The list price is $650. Here is how you get it: Let L=the list price L-0.3L=455 0.7L=455 L=455/0.7 L=$650

OpenStudy (mathhelp346):

but 650*.3 doesnt equal 455

OpenStudy (gw2011):

The net price is the discounted amount which means the list pric must be higher. Here is how it works: ($650)(0.3)=$195 which is the discount - this is subtracted from the list price to get the net as follows: $650-195=$455